OpenStudy (anonymous):

How would you simplify x^(4/3) over x^(5/6) and ^16 sqrt( x * x^3 * x^4 )

OpenStudy (anonymous):

For the first use the law \[{x^m \over x^n} = x^{m-n}\]

OpenStudy (jdoe0001):

\(\Large {\cfrac{1}{a^{\color{red} n}}=a^{-{\color{red} n}}\\ \quad \\ \quad \\ \cfrac{x^{\frac{4}{3}}}{x^{\frac{5}{6}}}\implies \cfrac{x^{\frac{4}{3}}}{1}\cdot \cfrac{1}{x^{\frac{5}{6}}}\implies \cfrac{x^{\frac{4}{3}}}{1}\cdot x^{-\frac{5}{6}}\implies x^{\frac{4}{3}}\cdot x^{-\frac{5}{6}}}\)

OpenStudy (jdoe0001):

\(\bf \Large {a^{\frac{{\color{blue} n}}{{\color{red} m}}} = \sqrt[{\color{red} m}]{a^{\color{blue} n}}\\ \quad \\ \quad \\ \sqrt[16]{x\cdot x^3\cdot x^4}}\)

OpenStudy (anonymous):

Which method is right? They dont result in the same answer.

OpenStudy (jdoe0001):

hmm what do you mean?